OpenStudy (anonymous):

Evan is making a rectangular flag in art class. He has created the main part of the flag, which measures 20 inches wide by 12 inches tall. He has exactly 228 square inches of material to create the flag’s border, and he wants to use all of it.

OpenStudy (anonymous):

pythagorean theorum is side1^2 plus side2^2 = longestside (#3) ^2 or a^2 + b^2 = c^2

OpenStudy (anonymous):

to get the border you need to find side c so 12*12 + 20*20 = 544 the square root of 544 is 23.32 and that is your 3rd side

OpenStudy (anonymous):

the area of a triangle is length times width times 1/2

OpenStudy (anonymous):

I'm sorry, I have to go.. :/ I hope this helps you start!

OpenStudy (anonymous):

The awnser is 23.32 20^2=400 12^2=144 400+144=544 The square root of 544 is 23.32

OpenStudy (anonymous):

since the sides are 12, 20 and 23.32 inches long (or 55.32 inches total) now you need to figure out how thick the border needs to be in order to use all of the cloth that you have.

OpenStudy (anonymous):

You will divide 228 by 55.32... or 228/55.32 you get just a little over 4 so, the border will be about 4 inches thick around your triangle.

OpenStudy (anonymous):

Sorry again for leaving yesterday, I hope you figured it out in time!